OpenStudy (anonymous):

Use the T-table below to determine 3 solutions to this equation: 4x - 2y = 8 x=0,2,-1 y=

OpenStudy (anonymous):

A quadratic equation is defined as an equation in which one or more of the terms ... In Geometry, we will concentrate on the graphical solutions to these systems. ... You can use the same table of values and simply find the y values for the straight line. ... Check (5,3) y = x2 - 4x - 2 3 = 52 - 4(5) - 2 3 = 3 check, y = x - 2 3 = 5 - 2